### v2

# Unsupervised Learning of Local Updates for Maximum Independent Set in Dynamic Graphs

Abstract
We present the first unsupervised learning model for finding Maximum Independent Sets (MaxIS) in dynamic graphs where edges change over time. Our method combines structural learning from graph neural networks (GNNs) with a learned distributed update mechanism that, given an edge addition or deletion event, modifies nodes’ internal memories and infers their MaxIS membership in a single, parallel step. We parameterize our model by the update mechanism’s radius and investigate the resulting performance–runtime tradeoffs for various dynamic graph topologies. We evaluate our model against a mixed integer programming solver and the state-of-the-art learning-based methods for MaxIS on static graphs (ICML 2020; NeurIPS 2020, 2023). Across synthetic and empirical dynamic graphs of 50–1,000 nodes, our model achieves competitive approximation ratios with excellent scalability; on large graphs, it significantly outperforms the state-of-the-art learning methods in solution quality, runtime, and memory usage. When generalizing to graphs of 10,000 nodes (100x larger than the ones used for training), our model produces MaxIS solutions 1.05–1.18x larger than any other learning method, even while maintaining competitive runtimes.
1 Introduction
Combinatorial optimization problems on graphs (e.g., Graph-Coloring, Traveling-Salesman, Vehicle-Routing, etc.) arise naturally in many practical domains (Applegate et al., 2007; Toth and Vigo, 2014). However, many of these problems are classically intractable to solve exactly, and some even resist efficient approximation. The situation is exacerbated by the fact that many real-world applications are dynamic, requiring algorithms to not only find solutions but also maintain them as the underlying structure evolves. While one could iteratively apply algorithms designed for static graphs at each time step, this approach is often computationally expensive and inefficient. Critically, this approach ignores the often close relationship between successive graph structures. A more effective strategy might avoid repeatedly recomputing solutions from scratch by instead leveraging these small structural changes to incrementally update an existing solution. Such update algorithms are an active area of theoretical algorithms research (Assadi et al., 2019; Hanauer et al., 2022; Chechik and Zhang, 2019; Behnezhad et al., 2019; Zheng et al., 2019; Gao et al., 2022), but no learning algorithms have yet taken this approach.
In this paper, we consider the Maximum-Independent-Set (MaxIS) problem in a dynamic setting. For static graphs, MaxIS is classically known to be NP-hard (Garey and Johnson, 1979) and hard to approximate: for general graphs, deterministic approximation within a constant factor is impossible (Robson, 1986) and within an $n^{1-\varepsilon}$ factor is NP-hard (Håstad, 1996). Nevertheless, MaxIS has a diverse range of real-world applications, from identifying functional nodes in brain networks (Afek et al., 2011) to constructing diversified investment portfolios (Hidaka et al., 2023). For example, different stocks can be modeled as nodes and those that are closely related have an edge between them to form a conflict graph; thus, the problem of building a diverse portfolio reduces to finding a high-value independent set. In many of these situations, the conflict graph changes dynamically over time due to factors such as neuronal development, stock market fluctuations, and other unpredictable events. Our goal is to solve MaxIS on such dynamically evolving graphs.
Recently, the learning community has developed several methods for solving combinatorial optimization problems on static graphs (Bengio et al., 2021). Central to many of these methods are Graph Neural Networks (GNNs) due to their efficiency in capturing the structural information present in real-world networks such as the Internet, social networks, and molecular interactions (Zhang et al., 2019; Wu et al., 2020). In the dynamic setting, recent models such as temporal GNNs are primarily concerned with tasks such as edge prediction, node classification, graph classification, etc. (Xu et al., 2020; Rossi et al., 2020). Work considering combinatorial optimization problems in a dynamic setting remains limited, with few examples such as Gunarathna et al. (2022), which is an indirect heuristic learning approach. No existing models learn update mechanisms comparable to theoretical, rule-based update algorithms.
We propose a model that solves MaxIS in dynamic graphs by directly learning an update mechanism instead of a heuristic. We utilize the power of message-passing GNNs (Battaglia et al., 2018) combined with sequence learning to learn an update mechanism that maintains an approximate MaxIS solution as the underlying graph changes. GNNs learn to aggregate structural information and update a node’s membership in the MaxIS solution at any instance of a dynamic graph, while sequence learning learns the effect of a change in the dynamic graph on the change in the MaxIS solution. Having learned an update mechanism, the model is able to update a MaxIS solution following a change in the dynamic graph in a single inference step, unlike more popular heuristic learning methods which require multiple inference steps and thus scale poorly with graph size. We underscore this as a necessary shift in perspective required to address some of the common criticisms of learning approaches to solving combinatorial optimization problems, namely, lack of scalability, computation inefficiency, and huge training sample requirements.
Contributions.
We present the first unsupervised learning model for solving Maximum-Independent-Set (MaxIS) on dynamic graphs. The central idea of learning local, distributed update mechanisms for combinatorial optimization problems on dynamic graphs is novel, and we demonstrate its efficiency in our evaluations. Compared to a commercial mixed integer programming solver (Gurobi Optimization, LLC, 2024), our model maintains favorable approximation ratios across a variety of dynamic graph topologies and sizes. For larger datasets, our model conclusively outperforms the three prior learning methods for MaxIS on static graphs: the current state-of-the-art, DP-GNN (Brusca et al., 2023); a single-shot method, Erdős-GNN (Karalias and Loukas, 2020); and a deep reinforcement learning method, LwDMIS (Ahn et al., 2020). Resource scaling for our method vs. DP-GNN is drastic; we obtain solutions up to 104x faster and use up to 27.5x less memory while producing up to 5.64x larger MaxIS solutions. Compared to Erdős-GNN and LwDMIS, our model delivers 1.071.22x larger MaxIS solutions than either method while also running 6.1824.8x faster and saving up to 2.7812.97x memory. These gains hold even in generalization experiments, where we evaluate our model on graphs 100x the size of those it was trained on.
2 Related Work
Traditional Algorithms for MaxIS.
Maximum-Independent-Set (MaxIS) is a classical NP-complete problem in graph theory and has been explored since the 1970s in the field of combinatorial optimization. Algorithms seeking exact solutions use reduction rules (Tarjan and Trojanowski, 1977), branching heuristics (Fomin and Kaski, 2013), and kernelization techniques (Lamm et al., 2016) to reduce the size of the search space. Among exact algorithms, the state-of-the-art is Xiao and Nagamochi (2017) which runs in $1.1996^{n}n^{O(1)}$ time and polynomial space, where $n$ is the number of nodes in the graph. Extensions of MaxIS to dynamic graphs remain relatively unexplored. Zheng et al. (2019) and Gao et al. (2022) have developed update algorithms using rule-based optimization techniques for incrementally maintaining an existing solution in response to edge insertions or deletions. However, no such approaches have been taken by the learning community, which we discuss next.
Learning in Dynamic Graphs.
Memory-based models utilize various forms of persistent, high-dimensional representations (i.e., memories) for nodes to learn the dynamics of a graph as it changes. These have been very successful in efficiently learning prediction and classification tasks in dynamic graphs (Rossi et al., 2020; Kumar et al., 2019; Trivedi et al., 2019; Yu et al., 2023). Some models also include a GNN-based graph embedding module to mitigate the problem of memory staleness, i.e., when a node’s memory hasn’t been updated for a long time. We utilize these innovations in our model. Other learning models for dynamic graphs do not involve memory (Xu et al., 2020; Pareja et al., 2020; Longa et al., 2023) but have never been applied to combinatorial optimization problems.
Learning for Combinatorial Optimization.
A vast body of machine learning research considers combinatorial optimization problems (Peng et al., 2021; Dai et al., 2017; Bengio et al., 2021). Among these, GNNs are the most common backbone given their excellent ability to characterize the structural information of graphs (Zhang et al., 2019; Wu et al., 2020). For MaxIS specifically, autoregressive approaches (i.e., those that take multiple inference steps to output a solution) such as reinforcement learning (RL) and dynamic programming (DP) have emerged as natural choices due to their easy constraint enforcement and a priori integral solution output (Brusca et al., 2023; Ahn et al., 2020). Yet, these methods suffer from sample inefficiency poor scalability, requiring a number of inference steps that scales linearly with problem size (Karalias and Loukas, 2020; Peng et al., 2021). Non-autoregressive approaches (i.e., "single-shot" methods that output a solution in one inference step) have also been proposed (Li et al., 2018; Karalias and Loukas, 2020; Wang et al., 2022; Schuetz et al., 2022), typically featuring unsupervised learning to avoid the issue of labeling a vast number of solutions. However, these techniques only investigate static graphs. To the best of our knowledge, the only other learning model that considers combinatorial optimization problems in dynamic graphs is by Gunarathna et al. (2022) applying RL to dynamic versions of the Traveling-Salesman and Vehicle-Routing problems. Our work is the first non-autoregressive, unsupervised learning method for a combinatorial optimization problem in dynamic graphs.
3 Maximum Independent Sets in Dynamic Graphs
Consider an undirected, static graph $G=(V,E)$ with nodes $V$ and edges $E$. An independent set $I\subseteq V$ of $G$ is a set of nodes no two of which are adjacent, i.e., for all $u,v\in I$, we have $(u,v)\not\in E$. Its size is its number of nodes, denoted $|I|$. A maximum independent set (MaxIS) of $G$ is any independent set $I$ of $G$ satisfying $|I|=\max\{|I^{\prime}|:I^{\prime}\text{ is an independent set of }G\}$.
Dynamic graphs (also called dynamic networks, temporal graphs, or evolving graphs) are graphs whose nodes and/or edges change over time (Bui-Xuan et al., 2003; Casteigts et al., 2012; Rossi et al., 2020; Kazemi et al., 2020). For this work, we consider dynamic graphs where at most one edge changes per time and the underlying set of nodes is fixed. Formally, a dynamic graph $\mathcal{G}$ is a finite sequence of graph snapshots $(G_{0},G_{1},\ldots,G_{T})$ where each snapshot $G_{t}=(V,E_{t})$ is an undirected, static graph on nodes $V$ and, for all time steps $0<t\leq T$, we have $|E_{t-1}\oplus E_{t}|=1$, i.e., there is exactly one edge addition or one edge deletion per time step. We denote the edge event at time $t$ which transitions $G_{t-1}$ to $G_{t}$ as $\mathcal{E}_{t}$. The neighborhood of a node $v$ at time $t$ is $\mathcal{N}_{t}(v)=\{u\in V\mid(u,v)\in E_{t}\}$ and its degree is denoted $d_{t}(v)=|\mathcal{N}_{t}(v)|$. The diameter of graph snapshot $G_{t}$, denoted $\text{diam}(G_{t})$, is the maximum hop-distance between any two nodes in $G_{t}$.
We define the Dynamic-Maximum-Independent-Set (Dynamic MaxIS) problem as follows: Given a dynamic graph $\mathcal{G}=(G_{0},G_{1},\ldots,G_{T})$—or, equivalently, an initial graph snapshot $G_{0}$ and a sequence of edge events $(\mathcal{E}_{1},\ldots,\mathcal{E}_{T})$—obtain a MaxIS for each snapshot $G_{t}$. This problem is clearly NP-hard since it contains the NP-complete MaxIS problem as a special case ($T=0$). Given a solution for the MaxIS problem, there exist a polynomial-time conversions to obtain solutions for the Minimum-Vertex-Cover and Maximum-Clique problems; hence, achieving good approximations for Dynamic MaxIS can also aid in approximating these problems’ dynamic versions.
4 An Unsupervised Learning Model for Dynamic MaxIS
As discussed in Section 2, memory-based models have been quite successful in recent years. Motivated by their effectiveness in predictive learning of dynamics, we present an unsupervised learning model focused on the MaxIS problem in dynamic graphs. We take inspiration from the Temporal Graph Networks (TGN) model proposed by Rossi et al. (2020) which generalizes memory-based models by including a GNN-based graph embedding module. We repurpose the two key modules of TGNs, memory and graph embedding, to learn an update mechanism that maintains a candidate MaxIS solution as the underlying graph changes. Moreover, inspired by the use of messages (seen as random perturbations) to improve GNN performance for NP-hard problems (Papp et al., 2021), we couple node memory to an event handling module to inform nodes about nearby edge events.
Figure 1: An overview of our unsupervised learning model for MaxIS on dynamic graphs. Nodes maintain memories (node colors) which are updated by signals propagating in an $\alpha$-radius from edge events (orange arrows). Each node within a $\beta$-radius of the edge event then locally aggregates its updated memory and those of its neighbors to estimate its new probability of MaxIS membership. These estimates are then used to compute loss (during training) or to generate an integral MaxIS solution (during testing and inference).
Our model responds to edge events (i.e., edge additions or deletions) over time (see Figure 1). At each time step, the event handling module is responsible for signaling all nodes within an $\alpha$ hop-distance of the edge event that the event occurred. These nodes then use the memory module to update their internal representations using this signal. Finally, each node within a $\beta$ hop-distance of the edge event uses the local aggregation (graph embedding) module to produce a probability of membership in the MaxIS solution (i.e., an estimate) based on its memory, its degree, and the memories of its neighbors. We describe each module and the training-related details of our model in the following subsections; pseudocode can be found in Appendix A.
Event Handling Module.
This module provides each node $v$ within an $\alpha$ hop-distance of the edge event $\mathcal{E}_{t}$ with a signal $s_{t}(v)$ encoding the type of event (addition or deletion) and the node’s normalized distance from the event. Formally, the signal for node $v$ at time $t$ is
$s_{t}(v)=[\texttt{enc}(\mathcal{E}_{t})~||~r_{t}(v)],$ (1)
where enc encodes edge deletions as $[0,1]$ and edge additions as $[1,0]$, and $r_{t}(v)$ is the hop-distance from $v$ to $\mathcal{E}_{t}$ normalized by $\alpha$ and linearly interpolated into [1,1][-1,1].
Memory Module.
A node’s memory is a high-dimensional representation maintained throughout the model’s execution that captures relevant structural changes affecting the node. At each time $t$, this module updates the memories of nodes within an $\alpha$ hop-distance of the edge event $\mathcal{E}_{t}$. Updates are performed using a GRU cell (Chung et al., 2014). Formally, the memory of a node $v$ at time $t$ is
[…]
Local Aggregation Module.
This module updates the estimates—i.e., probabilities of membership in the MaxIS solution—of all nodes within a $\beta$ hop-distance of the edge event $\mathcal{E}_{t}$. First, each node $v$ aggregates its immediate neighbors’ memories:
$\tilde{h}_{t}(v)=\texttt{ReLU}\left(\sum_{u\in\mathcal{N}_{t}(v)}\mathbf{W}_{1}m_{t}(u)\right),$ (3)
where $\mathbf{W}_{1}$ are learnable weights. Next, node $v$ combines these aggregated memories with its own memory and degree information to obtain a local embedding
[…]
where again $\mathbf{W}_{2}$ are learnable weights. Finally, node $v$ passes its embedding through a step-down MLP and a sigmoid function $\sigma$ to obtain its estimate (probability of MaxIS membership) in $[0,1]$:
$p_{t}(v)=\sigma\big{(}\texttt{MLP}(h_{t}(v))\big{)}.$ (5)
Any node $v$ beyond the $\beta$ hop-distance of $\mathcal{E}_{t}$ simply retains its previous estimate, i.e., $p_{t}(v)=p_{t-1}(v)$.
Update Training.
We first define a loss function from a node’s local perspective and then define a cumulative loss function from these local losses. Since the memory and local aggregation modules operate on individual nodes, a local loss function enables these modules to synergize and learn a distributed update mechanism. The loss for a node $v$ at time $t$ is
$\ell_{t}(v)=-p_{t}(v)+\frac{c}{2d_{t}(v)}\sum_{u\in\mathcal{N}_{t}(v)}p_{t}(u)p_{t}(v).$ (6)
The $p_{t}(v)$ term rewards nodes with large estimates (i.e., probabilities of MaxIS membership closer to 1), aligning with the goal of finding a large independent set. This is counterbalanced by the $\sum_{u\in\mathcal{N}_{t}(v)}p_{t}(u)p_{t}(v)$ term which penalizes violations of independence, i.e., when neighbors $u$ and $v$ both have large estimates. This constraint violation term is normalized by the degree of $v$ and is then halved to compensate for double-counting the loss from the perspectives of both $u$ and $v$. Finally, the constant $c$ is a hyperparameter balancing the two terms’ loss contributions; in practice, $c=3$ worked reasonably well across our evaluations (Section 5).
The cumulative loss at time $t$ is the sum of all local losses $\ell_{t}(v)$ for nodes $v$ within a $\beta$ hop-distance of the edge event $\mathcal{E}_{t}$ (i.e., all nodes that updated their estimates at time $t$). This set of nodes at this one time step forms a single batch in our training process. By setting $\beta$ as an appropriate function of dynamic graph size (see Model Variants, below), batch sizes will automatically scale to any input dynamic graph and provide sufficient inference samples for a training step. The sequential processing of all edge events in a dynamic graph’s training set forms a single epoch in our training process. In a single run, we train until the losses for epochs have stabilized and select the epoch with the least loss.
Model Variants.
As explained above, the hop-distances $\alpha$ and $\beta$ control which nodes utilize an edge event to update their memories and which nodes update their estimates and contribute to cumulative loss, respectively. In the bounded cascading (BCAS) model variant, we set $\alpha=\beta=\gamma\cdot\text{diam}(G_{0})$, where $\gamma\ll 1$, strictly bounding nodes’ memory and estimate updates to a local region around each edge event. This variant mimics rule-based approximate update algorithms for MaxIS which analogously spread out from the point of the edge event, updating nodes’ memberships to generate a new MaxIS (Behnezhad et al., 2019; Chechik and Zhang, 2019). Locally bounding these update regions typically yields very fast update operations, at the potential risk of missing better MaxIS solutions that require far-reaching updates.
In the non-cascading (No-CAS) model variant, we set $\alpha=0$ and $\beta=\text{diam}(G_{0})$; i.e., only the two nodes directly involved in the edge event update their memories, but all nodes in the graph update their estimates and contribute to cumulative loss. In this variant, the local aggregation module is primarily responsible for the graph-wide learning of the update mechanism, with node memory aiding in the locality of the edge event. This variant is important for dense graph topologies and edge dynamics where frequent update cascades may disrupt relevant information in nodes’ memories.
Initial MaxIS Generation and Unsupervised Learning.
Update algorithms for MaxIS crucially require an existing MaxIS solution for the initial snapshot $G_{0}$ which is later updated as the graph structure changes. We mitigate this (potentially very expensive) prerequisite by utilizing an initial MaxIS generation phase that efficiently produces memories for all nodes in $G_{0}$ and model weights that training can warm-start from. The idea is to build up $G_{0}$ one edge at a time (as a sequence of edge addition events), updating only the memories of the nodes involved in each edge addition. Once $G_{0}$ is fully constructed, estimates and corresponding losses are computed for all nodes. This constitutes a single epoch of this generation phase; in a single run of the generation phase, we execute epochs until the cumulative loss has stabilized. After executing multiple runs with random seeds, we extract the node memories and model weights from the run with the least cumulative loss to warm-start model training. Thus, the model as a whole learns to first find a candidate MaxIS (generation phase) and then maintain it (training phase) in an entirely unsupervised manner.
Integral Solution Generation.
During testing and deployment, we use a simple rounding and violation removal procedure to convert the model’s relaxed node estimates into a candidate MaxIS. Specifically, all nodes with estimates 0.5 or larger are initially included in the candidate solution; then, nodes violating independence are removed in decreasing order of number of violations until an independent set is obtained. Such a procedure is a necessary part of single-shot learning methods that directly use relaxation to solve optimization problems with hard constraints (Karalias and Loukas, 2020; Donti et al., 2021; Schuetz et al., 2022; Wang et al., 2022).
5 Experiments
In this section, we first describe our experimental methodology and then present and interpret the model evaluation results.
5.1 Experimental Methodology
Datasets.
We evaluate model performance across both synthetic and empirical dynamic graphs. For synthetic graphs, we consider Erdős–Rényi (ER) and Power Law (PL) graphs with randomly generated initial structures. For empirical graphs, we used initial structures given by GERMANY (Orlowski et al., 2010), a geographical network with a mesh topology; TWITTER (Leskovec and Krevl, 2014), a social network subgraph with a dense small-world topology; and BRAIN (Crossley et al., 2013), a biological network with a sparse small-world topology. Across all graphs, dynamics are produced by sampling edge additions or deletions according to the initial snapshot’s degree distribution for a desired number of time steps.
The dynamic graphs we consider span a range of sizes: small ($\approx$100 nodes, 50,000 time steps), medium ($\approx$1,000 nodes, 100,000 time steps), and large ($\approx$10,000 nodes, 5,000 time steps). Each dynamic graph is divided into its training, evaluation, and testing splits, arranged chronologically with increasing time steps. We use a 70:15:15 training, evaluation, testing split for small datasets and a 50:25:25 split for medium datasets. The large datasets only have 5,000 time steps because we use them solely for generalization experiments and do not train our model on them.
Our Model and Comparisons.
We ran our complete training pipeline using the two model variants: BCAS and No-CAS. For the BCAS variant, we set $\alpha=\beta=0.25\cdot\text{diam}(G_{0})$, i.e., half of the radius of the first snapshot of the dynamic graph. For each variant, we performed three training runs with different random seeds and report results for the median run. Since ours is the only learning method for MaxIS in the dynamic setting, we compare our method against a commercial mixed integer programming solver and existing learning methods for MaxIS on static graphs. Since these methods are intended for static graphs, we apply them to each snapshot of a dynamic graph independently.
  • Gurobi (Gurobi Optimization, LLC, 2024) is the most commonly used and commercially available mixed integer programming solver. We use Gurobi’s default settings for mixed integer program solving, which according to their documentation, is a branch-and-bound search.
  • DP-GNN (Brusca et al., 2023) is the state-of-the-art learning model for MaxIS in static graphs, combining dynamic programming with GNNs to form a multi-step, autoregressive comparator model.
  • Erdős-GNN (Karalias and Loukas, 2020) is an earlier GNN-based, single-shot, probabilistic learning model for combinatorial optimization in static graphs.
  • LwDMIS (Ahn et al., 2020) is a deep reinforcement learning method that adaptively adjusts its stages and defers decisions about nodes’ MaxIS membership to improve scalability.
All methods were trained (if applicable) and tested on a machine with an AMD EPYC 7413 CPU and a 20 GB slice of one NVIDIA A100 GPU. For the learning methods, training was halted after seven days. For Gurobi, we enforced a time limit of 1 s per snapshot; this is sufficient for producing exact solutions for small datasets, but yields approximations for medium and large ones.
Metrics.
We evaluate each method using three metrics: approximation ratio (performance), per-graph solution generation time (runtime), and peak memory usage. Performance refers to the mean ratio of a method’s candidate MaxIS size against that of Gurobi’s across all snapshots in a dynamic graph. Runtime refers to the mean time taken by a method to process a single snapshot and generate a candidate MaxIS across all snapshots in a dynamic graph (reported as seconds per graph, s/g). Peak memory usage denotes the maximum memory used by a method during training.
5.2 Results
(a) Performance
(b) Runtime
(c) Peak Memory Usage
Figure 2: Comparisons of models’ (a) performance, (b) runtime, and (c) peak memory usage across the seven small and medium datasets. The best method for each metric and dataset is marked with a ($\ast$); for performance, all methods for a given dataset whose mean performances are within one standard deviation (error bars) of the highest mean performance are considered “best”. Notably, DP-GNN is very slow on larger graphs, so its results for the medium datasets only reflect evaluation on the first 5,000 of 25,000 total testing timesteps. Also, we observed LwDMIS fail to train on some graph snapshots as its search space collapsed to null; thus, we only report peak training memory usage for those methods that successfully completed the entire training process.
The results highlight the efficient scalability and overall performance of our model. Both the BCAS and No-CAS variants of our model consistently outperform other learning-based methods in terms of runtime and peak memory usage, while also achieving comparable or superior performance across all graph sizes. Surprisingly, DP-GNN—although the current state-of-the-art—performs well only on small datasets. On the medium datasets, its performance is 1.30–4.20x worse than any next best method across datasets (Figure 2(a)) while running more than 1,000x slower (Figure 2(b)) and consuming up to 27.44x more memory (Figure 2(c)) than other methods. In contrast, earlier methods such as LwDMIS and Erdős-GNN exhibit better tradeoffs over different graph sizes: LwDMIS is reasonably fast but performs similarly or worse than Erdős-GNN, especially on small graphs, while Erdős-GNN achieves its stronger performance at the cost of slower speeds and worse scalability. Our model variants strike an effective balance between these, delivering 1.071.22x better performance than either method (Figure 2(a)) while also reducing runtime (6.1824.8x faster, Figure 2(b)) and saving up to 2.7812.97x memory (Figure 2(c)). This demonstrates the efficiency and scalability of our model when compared to any other existing learning method.
BCAS vs. No-CAS
Table 1: Pearson correlation for model performance and runtime vs. properties of graph snapshots. Correlation coefficients are calculated over all 75,000 graph snapshots in the testing splits of our medium datasets. Due to the large number of samples, we set our threshold for significance at $p<10^{-10}$. All reported correlations are significant with $p<10^{-40}$, except the one value marked with a ($\dagger$), which has $p=\text{0.00095}$.
Performance Runtime
Graph Properties BCAS No-CAS BCAS No-CAS
# Edges 0.99 0.97 –0.09 0.98
Min. Degree 0.91 0.94 –0.15 0.92
Max. Degree 0.96 0.89 0.12 0.92
Ave. Degree 0.99 0.97 –0.06 0.98
Ave. Clustering Coefficient 0.99 0.95 –0.01 0.97
Diameter –0.68 –0.79 0.43 –0.74
Ave. Node Distance –0.75 –0.85 0.39 –0.81
Table 1 shows correlations between our BCAS and No-CAS models’ performance and runtime vs. properties of the graph snapshots in the testing datasets. For performance, the same graph properties help and hurt both BCAS and No-CAS: both models perform better on graphs with many edges, high-degree nodes, and dense clustering but perform worse when diameter and node distance increase. This is intuitive, as both variants aggregate information from only their local neighbors and thus benefit from dense structure but suffer from distant changes that are not propagated. For runtime, there is a sharp contrast between the variants. BCAS runtimes are largely uncorrelated with graph properties—owing to its strictly local cascades and aggregation—exhibiting only a weak positive correlation with diameter and distance (stemming from the fact that $\alpha=\beta$ are set as a fixed fraction of diameter). No-CAS runtimes, on the other hand, are strongly determined by density (i.e., edge counts, degrees, and clustering), since every node in the graph must perform local aggregation at every time step and the complexity of local aggregation is entirely dependent on neighborhood size.
Generalization.
We highlight our model’s generalization ability by training our model variants on small synthetic datasets and then evaluating them on corresponding datasets 100x their size (Table 2). DP-GNN, Erdős-GNN, and LwDMIS display the same strengths and weaknesses as for the smaller datasets: DP-GNN scales so poorly that its runtime is four orders of magnitude worse than the others, Erdős-GNN achieves decent performance but is slower than the remaining methods, and LwDMIS runs very quickly at the cost of inferior performance. Our No-CAS variant, although not as fast as LwDMIS, achieves far superior performance to any other learning method and even ties Gurobi’s performance in a third of the runtime. On the PL dataset specifically, BCAS achieves the same great performance, but is 4.97x faster than No-CAS and 14.73x faster than Gurobi.
Table 2: Generalization results for large-sized datasets (10,000 nodes, 5,000 time steps). Performance is reported as the mean and standard deviation of MaxIS sizes across snapshots against those of Gurobi’s (run for 10 s). Runtime is reported in parentheses as mean seconds per graph snapshot. Notably, DP-GNN is prohibitively slow on these large graph snapshots, taking nearly a week just to evaluate the first six snapshots of each dataset. Thus, we report performance only for those methods that could be evaluated on the entire datasets.
Dataset ($\rightarrow$)
Method ($\downarrow$)
ER
($n=$ 10,000)
PL
($n=$ 10,000)
BCAS
0.63 $\pm$ 0.085
(0.015 s/g)
0.99 $\pm$ 0.002
(0.679 s/g)
No-CAS
1.17 $\pm$ 0.024
(5.349 s/g)
0.99 $\pm$ 0.001
(3.372 s/g)
DP-GNN
N/A
(36,239.508 s/g)
N/A
(39,994.408 s/g)
Erdős-GNN
1.07 $\pm$ 0.023
(19.202 s/g)
0.94 $\pm$ 0.009
(18.415 s/g)
LwDMIS
0.99 $\pm$ 0.020
(0.541 s/g)
0.92 $\pm$ 0.003
(0.296 s/g)
6 Conclusion
We present a single-shot, unsupervised learning model based on graph neural networks (GNNs) for the Maximum-Independent-Set (MaxIS) problem in dynamic graphs, where the topology evolves over time. Unlike heuristic learning approaches, our method learns an update mechanism analogous to those used in traditional rule-based algorithms for dynamic combinatorial optimization. We conduct experiments across various dynamic graph topologies, demonstrating that our model outperforms the DP-GNN (Brusca et al., 2023) and Erdős-GNN (Karalias and Loukas, 2020) models for MaxIS in terms of both runtime and memory scalability while maintaining strong approximation ratios even on dynamic graphs with up to 10,000 nodes. The prior LwDMIS model Ahn et al. (2020), although faster, lags far beyond our model’s performance. In the future, we plan to extend our model into a general framework for learning update algorithms applicable to a broader class of combinatorial optimization problems.
Limitations.
Our experiments demonstrate that our model outperforms existing learning methods for MaxIS in terms of approximation ratios, runtime, and training memory. However, a full comparison to traditional, non-learning heuristic methods is beyond the scope of this paper. Some preliminary experiments suggest that on large dynamic graphs, our model is faster than and achieves comparable performance to the iterative application of a greedy method—assuaging criticisms of learning-based approaches such as (Angelini and Ricci-Tersenghi, 2022)—but deterministic rule-based update algorithms that explicitly consider the dynamic setting may perform even better (Zheng et al., 2019; Gao et al., 2022). Also, the dynamics for all our datasets—synthetic or empirical—are sampled synthetically (though in a degree distribution-preserving way). It would be interesting to evaluate the performance of these methods on large dynamic graphs whose structure and dynamics are both empirical. Finally, our model currently makes limited use of temporal dynamics when learning its update mechanism, looking only at the effect of a single edge event. Future work is needed to leverage longer temporal dependencies and improve the update mechanism further.
Appendix A Model Pseudocode
Algorithms 13 show pseudocode for our model’s initial MaxIS generation procedure, update training process, and integral solution generation procedure, respectively. Throughout, graph snapshots $G_{t}$, edge events $\mathcal{E}_{t}$, signals $s_{t}(\cdot)$, node memories $m_{t}(\cdot)$, node estimates $p_{t}(\cdot)$, and node losses $\ell_{t}(\cdot)$ are defined as they were in the main text. Bold versions of these notations refer to the vector of all node properties of that type (e.g., $\mathbf{m}_{t}$ refers to all node memories). In a slight abuse of notation, we use $\mathcal{N}_{t}^{x}(\mathcal{E}_{t})$ to denote all nodes within an $x$ hop-distance from the edge event $\mathcal{E}_{t}$; i.e., the endpoints of the edge event are in $\mathcal{N}_{t}^{0}(\mathcal{E}_{t})$, those endpoints and their neighbors are in $\mathcal{N}_{t}^{1}(\mathcal{E}_{t})$, and so on.
Algorithm 1 Initial MaxIS Generation Epoch
1:Initialize all node memories as $\mathbf{m}_{0}\leftarrow\mathbf{0}$.
2:for all edges $(u,v)\in E_{0}$ do
3:  Let $\mathcal{E}$ be the edge event of adding $(u,v)$.
4:  Do event handling with signal $s\leftarrow[\texttt{enc}(\mathcal{E})~||~0]$ as in Eq. 1.
5:  Update $m_{0}(u)\leftarrow\texttt{GRU}([m_{0}(u)~||~s])$ and $m_{0}(v)\leftarrow\texttt{GRU}([m_{0}(v)~||~s])$ as in Eq. 2.
6:Compute node estimates $\mathbf{p}_{0}$ from the updated node memories $\mathbf{m}_{0}$ as in Eqs. 35.
7:Compute node losses $\mathbf{\ell}_{0}$ from $\mathbf{p}_{0}$ as in Eq. 6.
8:Compute cumulative loss $L_{0}\leftarrow\sum_{v\in V}\ell_{0}(v)$.
Algorithm 2 Update Training Epoch
1:Initialize node memories $\mathbf{m}_{0}$ using Algorithm 1.
2:for all edge events $\mathcal{E}_{t}\in(\mathcal{E}_{1},\mathcal{E}_{2},\ldots,\mathcal{E}_{T})$ in the training set do
3:  for all nodes $v\in\mathcal{N}_{t}^{\alpha}(\mathcal{E}_{t})$ do
4:   Do event handling with $s_{t}(v)\leftarrow[\texttt{enc}(\mathcal{E}_{t})~||~r_{t}(v)]$ as in Eq. 1.
5:   Update $m_{t}(v)$ using $s_{t}(v)$ as in Eq. 2.
6:  for all nodes $v\in\mathcal{N}_{t}^{\beta}(\mathcal{E}_{t})$ do
7:   Compute estimate $p_{t}(v)$ from $m_{t}(v)$ as in Eqs. 35.
8:  for all nodes $v\in\mathcal{N}_{t}^{\beta}(\mathcal{E}_{t})$ do
9:   Compute node loss $\ell_{t}(v)$ from $\mathbf{p}_{t}$ as in Eq. 6.
10:  Compute cumulative loss $L_{t}\leftarrow\sum_{v\in\mathcal{N}_{t}^{\beta}(\mathcal{E}_{t})}\ell_{t}(v)$.
Algorithm 3 Integral Solution Generation (Testing & Inference)
1:Initialize node memories $\mathbf{m}_{0}$ using Algorithm 1.
2:for all edge events $\mathcal{E}_{t}\in(\mathcal{E}_{1},\mathcal{E}_{2},\ldots,\mathcal{E}_{T})$ in the testing set do
3:  for all nodes $v\in\mathcal{N}_{t}^{\alpha}(\mathcal{E}_{t})$ do
4:   Do event handling with $s_{t}(v)\leftarrow[\texttt{enc}(\mathcal{E}_{t})~||~r_{t}(v)]$ as in Eq. 1.
5:   Update $m_{t}(v)$ using $s_{t}(v)$ as in Eq. 2.
6:  for all nodes $v\in\mathcal{N}_{t}^{\beta}(\mathcal{E}_{t})$ do
7:   Compute estimate $p_{t}(v)$ from $m_{t}(v)$ as in Eqs. 35.
8:  for all nodes $v\in\mathcal{N}_{t}^{\beta}(\mathcal{E}_{t})$ do $\triangleright$ Estimate Rounding
9:   if $p_{t}(v)\geq 0.5$ then $I_{t}(v)\leftarrow 1$.
10:   else $I_{t}(v)\leftarrow 0$.
11:  while there exists a node $v\in\mathcal{N}_{t}^{\beta}(\mathcal{E}_{t})$ with violations do $\triangleright$ Violation Removal
12:   Choose the node $v\in\mathcal{N}_{t}^{\beta}(\mathcal{E}_{t})$ with the most violations.
13:   Remove $v$ from the MaxIS wtih $I_{t}(v)\leftarrow 0$.